OpenStudy (anonymous):

Find y when x=24 Y varies directly as x When y=30, x=9.

OpenStudy (anonymous):

7.2???

OpenStudy (aravindg):

y varies directly with x ie y=kx

OpenStudy (aravindg):

Now when y=30 x=9

OpenStudy (aravindg):

Substitute in equation and find k

OpenStudy (aravindg):

Once you get k ,you are given x=24 .Find y .

OpenStudy (anonymous):

So how would the equation look in order to find k?

OpenStudy (aravindg):

30=k *9

OpenStudy (anonymous):

Ok, i am at a lost on how to find k from that, i'm not just asking you to give me the answer i actually want to learn how to do solve this.

OpenStudy (anonymous):

hmm your lost but you want the answer?

OpenStudy (aravindg):

30= k X 9 k=30/9 .Was that really hard?

OpenStudy (aravindg):

I hope you can do the rest by yourself.

OpenStudy (anonymous):

So i got 3.33

OpenStudy (anonymous):

y=kx When y=30, x=9. 30=k *9 divide by 9 on both sides to isolate k 30/9 = k 3.333 = k ----------------------------------- Find y when x=24 y = kx y = 3.333 * 24 y = 80

OpenStudy (anonymous):

basically what AravinG said :)

OpenStudy (anonymous):

Oh wow that made it very simple, thank you very much. I'll go ahead and apply that to the original question :)

OpenStudy (anonymous):

yes that is a great idea :) any doubts or questions feel free to ask :)